In an article in the Journal of Advertising, Weinberger and Spotts compare the use of humor in television ads in the United States and the United Kingdom. They found that a substantially greater percentage of U.K. ads use humor. (a) Suppose that a random sample of 380 television ads in the United Kingdom reveals that 140 of these ads use humor. Find a point estimate of and a 95 percent confidence interval for the proportion of all U.K. television ads that use humor. (Round your answers to 3 decimal places.) pˆ = The 95 percent confidence interval is [ , ]. (b) Suppose a random sample of 470 television ads in the United States reveals that 112 of these ads use humor. Find a point estimate of and a 95 percent confidence interval for the proportion of all U.S. television ads that use humor. (Round your answers to 3 decimal places.) pˆ = The 95 percent confidence interval is [ , ]. (c) Do the confidence intervals you computed in parts a and b suggest that a greater percentage of U.K. ads use humor? , the U.K. 95 percent confidence interval is the maximum value in the confidence interval for the U.S.
Answer #1

a) \widehat p = 140/380 = 0.368

At 95% confidence level, the critical value is z0.025 = 1.96

The 95% confidence interval is

\widehat p +/- z0.025 * sqrt(\widehat p(1 - \widehat p )/n)

= 0.368 +/- 1.96 * sqrt(0.368 * (1 - 0.368)/380)

= 0.368 +/- 0.048

= 0.320, 0.416

b) \widehat p = 112/470 = 0.238

The 95% confidence interval is

\widehat p +/- z0.025 * sqrt(\widehat p(1 - \widehat p )/n)

= 0.238 +/- 1.96 * sqrt(0.238 * (1 - 0.238)/470)

= 0.238 +/- 0.039

= 0.199, 0.277

c) Since the interval width in part (a) is wider than the interval width in part(b) , so the intervals suggests that a greater percentage of U.K. ads use humor.